A sample of oxygen gas at 20.0°C has a volume of 26.0L and exerts a pressure of 640.0mmHg. What is the mass of this gas?


1
Expert's answer
2021-04-02T05:13:10-0400

"T = 20 + 273 = 293 \;K \\

V = 26.0 \;L \\

p = 640.0 \; mmHg = 0.8421 \;atm"

Ideal gas law

pV = nRT

R= 0.08206 L×atm/mol×K

"n = \frac{pV}{RT} \\

= \frac{0.8421 \times 26.0}{0.08206 \times 293} \\

= 0.91 \;mol \\

m = n \times M"

M(O2) = 32 g/mol

"m = 0.91 \times 32 = 29.14 \;g"

Answer: 29.14 g